OpenStudy (anonymous):

WILL MEDAL AND FAN ALL WHO HELP What is the solution to the equation? -4x + 5(x-2)= -16 + 10

OpenStudy (anonymous):

x=-16

OpenStudy (ankit042):

-4x+5x-10=-16+10 x-10=-6 solve now

OpenStudy (anonymous):

witch one is it

OpenStudy (anonymous):

x=-16

OpenStudy (anonymous):

how do i know if @ankit042 is right ?

OpenStudy (anonymous):

check

OpenStudy (ankit042):

there is actually a way to find it out put the value of x in our original equation and left hand side should be equal to right hand side

OpenStudy (ankit042):

and Yes I know I am right

OpenStudy (anonymous):

ight ima see whos right brb

OpenStudy (anonymous):

i got 4 @ankit042 and @hakerman

OpenStudy (ankit042):

see I was correct !

OpenStudy (anonymous):

if you got the answer way do you need us

OpenStudy (anonymous):

u got 4 too ? @ankit042 and hankrman howd u get -16?

OpenStudy (lolakammy5o1o6):

Bill what do you think it is?

OpenStudy (anonymous):

i use a calcelater

OpenStudy (anonymous):

o i did it again and got 4

OpenStudy (anonymous):

http://prntscr.com/53np0a